5. Geoffrey’s Toy Box currently has credit sales of $1,000,000. Sixty percent of customers pay on day 20, 30 percent pay on day 40 and 10 percent pay on day 60. The new CFO estimates that offering longer credit terms would (1) increase the days sales outstanding to 50 days and (2) increase sales to $1,200,000. However, bad debt losses, which were 2 percent on the old sales, would increase to 5 percent on the incremental sales while bad debts on the current level of sales would stay at 2 percent. Variable costs are 78 percent of sales, and Geoffrey’s Toy Box has a 16 percent line of credit. Given corporate taxes of 21%, what would the annual incremental after-tax profit be if Geoffrey’s Toy Box extended its credit period? (10 points)
Answer #1
The costs and benefits associated with the proposal to extend credit are calculated below:
1] Incremental contribution = (1200000-1000000)*(1-78%) = $            44,000
2] Incremental bad debts = -200000*5% = $         -10,000
3] Days' sales outstanding [current] = 20*60%*40+30%+60*10% = 30
Investment in receivables [current] = 1000000*(30/365)*78% = $           64,110
Investment in receivables [proposed] = 1200000*(50/365)*78% = $      1,28,219
Increase in investment [128219-64110] $           64,110
Interest expense on additional investment = 64110*16% = $           -10,258
Annual incremental before tax profit $            23,742
Tax at 21% $              4,986
Annual incremental after tax profit $            18,756